If 9 and x have the same average as 3,7, and 14 what is the value of x?
Answer by stanbon(57206)

If 9 and x have the same average as 3,7, and 14 what is the value of x?
----
Average of 9 and x: (9+x)/2
----
Average of 3,7,14: (3+7+14)/3 = 8
---
Solve (9+x)/2 = 8
9+x = 16
x = 7
===========
Cheers,
Stan H.
===========